OpenStudy (anonymous):

Evaluate the definite integral:

OpenStudy (anonymous):

\[\int\limits_{4}^{7}t^3\ln(2t)dt\]

OpenStudy (anonymous):

If I did integration by parts would the 4 and 7 be in the integral left over?

OpenStudy (mertsj):

Wolf says 1316.95

OpenStudy (anonymous):

Ok.... how did you get that?

OpenStudy (mertsj):

1/16t^4(4log(2t)-1)

OpenStudy (turingtest):

integration by parts\[u=\ln(2t)\]\[dv=t^3\]\[\int t^3\ln(2t)dt=\frac1 4t^4\ln(2t)-\frac1 4\int t^3dt=\frac1 4t^4\ln(2t)-\frac{1}{16}t^4\]\[=\frac{1}{16}t^4(4\ln(2t)-1)\]evaluate from 4 to 7 forgot parentheses earlier